what equation would shift the following parabola to the right 5 units f(x)=-7x^2+5?: what equation would shift the following parabola to the right 5 units f(x)=-7x^2+5?
Answer by jim_thompson5910(9869)

If you replace x with x-5, then the entire parabola will shift 5 units.


f(x)=-7x^2+5 Start with the given equation


f(x-5)=-7(x-5)^2+5 Replace x with x-5


f(x-5)=-7(x^2-10x+25)+5 FOIL


f(x-5)=-7x^2+70x-175+5 Distribute


f(x-5)=-7x^2+70x-170 Combine like terms


So the new equation (call it g(x)) that is shifted over 5 units to the right is

g(x)=-7x^2+70x-170